40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4876     Entry time: Fri Jun 24 07:40:23 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: PRM damping restored 

The PRM sus damping restored. C1:SUS-PRM_SDPD_VAR is still 20-30mV and going up.  Side gain  turned on. This pulled it down to 5-8 mV

Why is the side osem sensing voltage 4.4V ? It can not be higher than ~2.4V.......something is rotten in the state of Denmark?

Edit by KI:

 It's because Valera increased the transimpedance gain of the PRM SIDE OSEM to match the signal level to the new ADC range (#3913 ).

ELOG V3.1.3-